compared to a single lamp connected to the battery two identical lamps
Answer #1

Answer is Less current

Current flowing when single lamp connected to a battery

I1=E/R

Current flowing when two lamps connected in series to a battery

I2=E/(R+R)=(1/2)(E/R)=(1/2)I1